Can someone please help, I put the problem down then I wrote it out in word form just in case it didn't make since in problem form. I do not understand how to solve for x, thanks in advance.
Solve each of the following for x, where x is a real number.
1. x√3-2=5x√3 wrote out: (x square root 3 minus 2)equals(5x square root 3)
2. 0.9-x=1 (0.9=.99..) wrote out: 0.9 the nine has a line over the top of it -x=1 ( 0.9 ( nine has a line over top it = .99.....)
3. 0.4 + x=1 (0.4=.44....)wrote out: 0.4 ( line over top 4)+x=1 ( 0.4 (line over top 4)=.44....)
4. 5.2x-0.01<0.2x+3.6

Answer by stanbon(75887)

Solve each of the following for x, where x is a real number.
1. x√3-2=5x√3 wrote out: (x square root 3 minus 2)equals(5x square root 3)
Subtract xsqrt(3) from both sides to get:
-2 = 4xsqrt(3)
xsqrt(3) = -1/2
Multiply both sides by sqrt(3) to get:
3x = (-1/2)sqrt(3)
x = (-1/6)sqrt(3)
---------------------------------

2. 0.9-x=1 (0.9=.99..) wrote out: 0.9 the nine has a line over the top of it -x=1 ( 0.9 ( nine has a line over top it = .99.....)
---
0.99999 = 1 (this can be shown)
Your equation: 1 - x = 1
x = 0
----------------------------------

3. 0.4 + x=1 (0.4=.44....)wrote out: 0.4 ( line over top 4)+x=1 ( 0.4 (line over top 4)=.44....)
----
x = 1-0.4444444
x = 0.555555555555...
----------------------------------
4. 5.2x - 0.01 < 0.2x + 3.6
5x < 3.61
x < 0.725
==================================
Cheers,
Stan H.